A body of mass 4.7 kg makes an elastic collision with another body at rest and continues to move in the original direction but with 1/3 of its original speed. (a) What is the mass of the other body? (b) What is the speed of the two-body center of mass if the initial speed of the 4.7 kg body was 6.6 m/s?
Answer #1

We know that the energy of system remains conserved

So

By Applying Law of conservaton of energy

Let the mass of second body is "m"

And let the speed of mass 4.7 kg be "v"

(a)

By the conservation of kinetic enegy

4.7 * v^{2} = (4.7+m)*(v/3)^{2}

By solving this equation

m = 37.6 kg

(b)

If the speed of 4.7 kg = 6.6 m/s

then as given in question after collision speed will be 1/3 So

Speed of centre of mass of two bodies = 6.6/3 m/s

= 2.2 m/s